# Minimax Decision Trees via Martingale Approximations

### v1

@@ 3.2 Cyclic MinimaxSplit risk bound
Remark 8.
Similar considerations also apply for Theorems 9 and 10 below.
Let us explain the intuition behind Theorem 6 when $d=1$ and $Y=g(X)$ for some $g\in\mathcal{G}$. Roughly speaking, (24) amounts to
$\mathbb{E}[(Y-M_{k})^{2}]\leqslant L2^{-2k/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}.$ (26)
To verify (26), observe that $\{M_{k}\}_{k\geqslant 0}$ is a martingale converging to $Y$. We employ the tree representation to control $\mathbb{E}[(M_{k}-M_{k+1})^{2}]$. By Hölder’s inequality and (21),
$\mathbb{E}[(M_{k}-M_{k+1})^{2}]$ $=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}$
$\leqslant\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}\bigg{)}^{1/3}$
×$\hskip 28.45274pt\times\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}$
$\leqslant\mathrm{Var}(Y)^{2/3}2^{-2k/3}\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}.$
$\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\leqslant\sum_{A\in\pi_{k+1}}\Big{(}\sup_{A}g-\inf_{A}g\Big{)}\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}.$
On the other hand, $\mathrm{Var}(Y)=\mathrm{Var}(g(\mathbf{X}))\leqslant\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$. Altogether (and using the martingale property), we obtain $\mathbb{E}[(Y-M_{k})^{2}]\leqslant L2^{-2k/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}$.
Let us compare the rates between (23) and (24). For $Y=g(\mathbf{X})$ with $g\in\mathcal{G}$ and $d$ fixed, we consider $\varepsilon>0$. For the VarianceSplit algorithm, (23) requires $2^{k}\approx e^{L\varepsilon^{-1}}$ nodes in the decision tree to guarantee that the MSE is smaller than $\varepsilon$ for some $L>0$. On the other hand, for the cyclic MinimaxSplit algorithm, (24) requires only $2^{k}\approx\varepsilon^{-Ld}$ nodes. Therefore, as $\varepsilon\to 0$, the cyclic MinimaxSplit algorithm requires fewer nodes to attain the same risk bound and thus facilitates computation.
Note that Theorem 6 exhibits the curse of dimensionality, which cannot be circumvented in the current setting to the best of our knowledge (Cattaneo et al., 2022; Tan et al., 2022).
3.3 Empirical risk bound for cyclic MinimaxSplit
We now focus on the regression setting and apply Theorem 9 to derive finite-sample performance guarantees. Suppose that under the original law $\mathbb{P}_{*}$, $\mathbf{X}_{*}$ is marginally atomless and
$Y_{*}\mid\mathbf{X}_{*}=g_{*}(\mathbf{X}_{*})+\varepsilon:=\mathbb{E}[Y_{*}\mid\mathbf{X}_{*}]+\varepsilon,$
where the error $\varepsilon=Y_{*}-g_{*}(\mathbf{X}_{*})$ is sub-Gaussian, i.e., for some $\sigma>0$,
$\mathbb{P}_{*}(|\varepsilon|\geqslant u)\leqslant 2\exp\Big{(}-\frac{u^{2}}{2\sigma^{2}}\Big{)},\quad u\geqslant 0.$
Also, we consider the empirical law of $(\mathbf{X},Y)$ of $N$ samples from $\mathbb{P}_{*}$.
A limitation of Theorem 6 is that it applies only if the covariate $\mathbf{X}$ is marginally atomless (for (20) to hold). Therefore, we need the following result that deals with measures whose marginals may contain atoms. This also provides a more complete answer to (P-general).
Theorem 9.
Suppose that $\mathbf{X}$ is purely atomic [^4] The atomicity condition is only required such that the $\operatorname*{arg\,min}$ in (15) is well-defined; see Footnote 3. and for some $N>0$,
Let $M:=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$ and the coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$ be constructed from the cyclic MinimaxSplit algorithm. Then uniformly for any $\delta>0$ and $k\geqslant 0$,
$\begin{split}\mathbb{E}[(Y-M_{k})^{2}]&\leqslant(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}+\inf_{g\in\mathcal{G}}\bigg{(}\Big{(}(1+\delta)+\frac{2(1+\delta)(1+\delta^{-1})}{3\cdot 2^{2\lfloor k/d\rfloor/3}}\Big{)}\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}]\\ &\hskip 170.71652pt+(1+\delta^{-1})\bigg{(}\frac{2}{3}+\Big{(}\frac{1+\delta^{-1}}{4}\Big{)}^{2/3}\bigg{)}2^{-2\lfloor k/d\rfloor/3}\left\lVert g\right\rVert_{\mathrm{TV}}^{2}\bigg{)}.\end{split}$ (28)
The compensation term
$(1+\delta^{-1})2^{-2\lfloor k/d\rfloor/3}\frac{2^{k+2}M^{2}}{3N}$
appearing in (28) guarantees that if the number of samples $N$ is much larger than $2^{k}M^{2}$, the asymptotic upper bounds (24) and (28) are comparable.
[…]
Theorem 10.
3.4 Numerical experiments
In this section, we provide experiments that verify the advantages of MinimaxSplit and cyclic MinimaxSplit methods in one- and two-dimensional domains, as well as integrated ensemble methods applied to approximating a real-valued function, and therefore it can be used to the application of denoising images. Further numerics (including those in high dimensions) are provided in Appendix B.
3.4.1 One-dimensional input domain
Figure 1: The comparison of the optimal tree, Scikit-learn DecisionTreeRegressor, our MinimaxSplit tree, and VarianceSplit tree on a sinusoidal target function $y=f(x)+\varepsilon=\sin(4\pi x)+\varepsilon$, $x\in[0,1],\varepsilon\sim N(0,\sigma_{\varepsilon}^{2})$ with low (first row, noise variance $\sigma_{\varepsilon}^{2}=0.01$) and high (second row, noise variance $\sigma_{\varepsilon}^{2}=0.25$) noise levels. We illustrate the true function $f$ using solid red curves and use the same training set for each of the low-noise and high-noise cases; the training set represented by the solid dots is one of the 10 different training sets of size 100. We display 10 different model fits based on 10 different batches of training sets, and the averaged mean-squared error evaluated on a different training set across 10 batches.
As shown in Figure 1, our first experiment compares decision tree regression methods under varying noise conditions and offers notable insights into their relative performance and characteristics. The optimal tree is obtained using global optimization over all possible split values when conditioning on the tree structure and split coordinates (in our case, split coordinate can only be chosen along one coordinate since $x\in\mathbb{R}$). Specifically, for max_depth=3, it employs a fixed structure with $72^{3}-1=$ 7 split points, optimized simultaneously using sequential least squares programming to minimize the overall MSE on the training data. The topology of the tree is predetermined and enforced through constraints in the optimization process. The optimal tree approach contrasts with traditional variance-based, top-down decision tree algorithms by optimizing all splits at once, potentially capturing global data patterns more effectively. However, its fixed structure may limit adaptability to varying data complexities compared to conventional adaptive tree methods. The optimal tree serves as a baseline, particularly in high-noise scenarios, and suggests that its optimization criteria may be more effective at capturing the underlying function while minimizing overfitting.
On the other hand, our MinimaxSplit’s competitive performance, especially under high noise, aligns with the theoretical strengths of the minimax optimization in decision trees. Its ability to maintain fidelity to the underlying function in noisy conditions underscores the value of robust splitting criteria in challenging environments. This observation may motivate further exploration of minimax-based approaches in domains with uncertain data quality or high noise levels. As another reference, the standard Scikit-learn DecisionTreeRegressor and the decision tree splitting based on variance minimization show similar performance patterns, with notable degradation in high-noise scenarios. This similarity is expected given their likely use of variance reduction as the splitting criterion. However, their performance lags behind that of MinimaxSplit in noisy conditions, highlighting the flexibility of MinimaxSplit in the presence of significant noise. The varying relative performance suggests that ensemble methods leveraging the strengths of different approaches could yield more robust and adaptive models across diverse data conditions, which we detail in Appendix B. In conclusion, this study provides valuable insight into the behavior of decision tree regression methods under varying noise conditions, highlighting the potential for improved performance through custom optimization criteria and robust splitting methods.
3.4.2 Two-dimensional input domain
For this experiment, we use a synthetic dataset generated from a complex, non-linear function of two features, $\mathbf{x}=(x_{1},x_{2})$. The function is designed to simulate realistic challenges encountered in regression tasks, such as non-linearity and interaction between features. Specifically, the true function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ used for generating the target values is defined as:
$\begin{split}f(\mathbf{x})&=\left(2-2.1u_{1}^{2}+\frac{u_{1}^{4}}{3}\right)u_{1}^{2}+u_{1}u_{2}+\left(-4+4u_{2}^{2}\right)u_{2}^{2};\\ &u_{1}=3(x_{1}-0.5),\\ &u_{2}=3(x_{2}-0.5),\end{split}$ (30)
where $x_{1},x_{2}$ are independently and uniformly sampled from $[0,1]$. This function incorporates both quadratic and higher-order polynomial terms, creating a complex surface with multiple peaks and valleys. The visualizations of the prediction surfaces, as shown in Figure 2, further illustrate the improved accuracy and smoother transitions in the predicted values when using the proposed methods.
Figure 2: The predictions and true function values are visualized using heatmaps, allowing for a clear comparison of how well each model approximates the underlying function (30). The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models. The next three subplots depict the predictions of the $L^{1}$ VarianceSplit, $L^{1}$ MinimaxSplit, $L^{2}$ VarianceSplit, and $L^{2}$ MinimaxSplit. Top Row: max depth 2; Middle Row: max depth 6; Bottom Row: max depth 10.
In this example (as well as in Sections 3.4.3 and 3.4.4 below), we also incorporate an $L^{1}$ variation of the VarianceSplit and MinimaxSplit algorithms, where in the splitting rule, we replace the variances in (11), (14), and (15) by the $L^{1}$ distance from the mean. For instance, the $L^{1}$ variation of (11) is
$\begin{split}(j,x_{j})&=\operatorname*{arg\,min}_{(j,\,x_{j})}\Big{(}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathbb{E}\big{[}|Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}]|\mid\mathbf{X}\in A,\,X_{j}<x_{j}\big{]}\\ &\hskip 56.9055pt+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathbb{E}\big{[}|Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}]|\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\big{]}\Big{)}.\end{split}$ (31)
The optimal $L^{1}$ split point may not be unique, in which case we break ties arbitrarily.
3.4.3 Application to denoising images
Another notable experiment involves applying decision trees to denoise noisy images (Luo et al., 2024). We apply the different tree variants to predict pixel values based on their locations, effectively treating this as a regression problem. The experiments include the following configurations:
  1. 1.
    VarianceSplit with $L^{1}$ (see (31)) and $L^{2}$ (see (11)) norms: These trees use either $L^{1}$ or $L^{2}$ norms to measure errors. Variance $L^{1}$ trees tend to capture more outlier noise, while $L^{2}$ norm-based trees provide smoother approximations.
  2. 2.
    MinimaxSplit and cyclic MinimaxSplit: These methods, especially when paired with cyclic padding, are effective in balancing feature splits in high-dimensional image data. The cyclic MinimaxSplit method shows robustness against feature dominance, which often occurred when a few features had much stronger correlations than others.
The results show that mixed methods, where the error calculation alternates between minimax and variance-based splits, lead to a balance between local feature capturing and overall smoothness.
Figure 3: The noiseless image, noisy image and denoised images using different kinds of splitting criteria for a single tree, allowing for a detailed comparison of how well each model approximates the geometric structure. The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models. The next four subplots depict the predictions of the VarianceSplit, Weighted VarianceSplit, MinimaxSplit, cyclic MinimaxSplit for $L^{1}$ (top row) and $L^{2}$ (bottom row) decision trees. All trees are fitted with a max depth of 10.
The experiment shown in Figure 3 compares decision tree regression methods for denoising the classic Astronaut (Van der Walt et al., 2014) grayscale image, which has been preprocessed to 128 by 128. Subsequently, Gaussian noise with a standard deviation of 0.25 was added to create the noisy version. Examining the output plot reveals distinct characteristics of each method. The $L^{1}$ VarianceSplit approach (RMSE=0.140421) produces a piecewise constant approximation with visible rectangular artifacts, indicating a tendency to oversegment the image. The $L^{1}$ MinimaxSplit method (RMSE=0.114668) yields a smoother reconstruction with better preservation of large-scale features and edges, particularly evident in the feather of the subject’s outline. The $L^{2}$ VarianceSplit method (RMSE=0.138452) results in a slightly smoother output compared to its $L^{1}$ counterpart, but still exhibits noticeable blockiness. The $L^{2}$ MinimaxSplit approach (RMSE=0.113193) achieves the lowest RMSE, producing a reconstruction that balances smoothness with feature preservation, especially apparent in the gradual shading of the astronaut’s shoulder and the nuanced details around her head. The superior performance of the MinimaxSplit criteria, especially with the $L^{2}$ norm, can be attributed to its focus on minimizing worst-case errors. This strategy is particularly effective for image denoising, where it helps preserve important structural elements while smoothing out noise.
Unlike traditional methods that select the best feature at each split based solely on immediate error reduction, the cyclic MinimaxSplit method systematically cycles through the available features as the tree grows. At different levels of the tree, different features are chosen for splitting, based on a pre-determined order or cyclic pattern. This ensures that all features are considered fairly and thus reduces the risk of the model becoming overly dependent on a small subset of features. The cyclic nature emphasizes on modeling each dimension of the underlying function while avoiding heterogeneity across dimensions, which can be particularly beneficial in multivariate cases where feature correlations are complex.
3.4.4 From single tree to ensemble
The ensemble method we consider here is the random forest (Breiman et al., 1984), which aggregates multiple decision trees to improve the precision and robustness of the prediction. Each tree is grown to a fixed depth $k$, and different combinations of splitting criteria are used to determine how each subset of the data is partitioned. The ensemble predictions are then averaged to provide the final output (see Algorithm 1 in Appendix C).
Specifically, we use the MinimaxSplit, cyclic MinimaxSplit, and VarianceSplit algorithms as the base learners in a random forest setting. Each tree in the ensemble is trained on a bootstrap sample of the original dataset and predictions are made by averaging the output of individual trees, effectively reducing the overall variance of the model. The model aims to maintain a balance between bias and variance by taking advantage of the strengths of each base model while averaging individual weaknesses.
$L^{2}$ norm Refer to caption
$L^{1}$ norm Refer to caption
Figure 4: Predictions and true function values of (30), visualized using heatmaps. Top Row: The first subplot shows the heatmap of the true function values across the input space, serving as the benchmark for evaluating the models. The next three subplots depict the predictions of the variance-based decision tree, minimax decision tree, and cyclic minimax decision tree (depth $k=10$ with $L^{1}$ (top) and $L^{2}$ (bottom) norm), respectively. Each plot is annotated with the corresponding MSE to quantitatively assess the accuracy. Bottom Row: The final four subplots visualize the predictions from four different random forest models (using Algorithm 1 of Appendix C), each built with varying combinations of the three error methods. Again, each plot is labeled with the MSE to facilitate direct comparison.
We apply the above ensemble approach in the same setting as in Section 3.4.2. The experiment in Figure 4 involves training multiple regression models (VarianceSplit, MinimaxSplit, and cyclic MinimaxSplit) and ensemble models on the generated dataset. The results demonstrate that the MinimaxSplit+MinimaxSplit ensemble outperforms the traditional variance-based tree in terms of MSE, particularly in scenarios where the data distribution is non-uniform. Further discussions on a weighted aggregation regime will be provided in Appendix B.3.
4 Contribution and Future Work
In this study, we introduced a novel decision tree splitting strategy, the MinimaxSplit algorithm, along with its multivariate variant, the cyclic MinimaxSplit algorithm. Unlike traditional VarianceSplit methods, which aim to reduce overall variance in the children nodes, our MinimaxSplit algorithm seeks to minimize the maximum variance within the split partitions, thereby reducing the risk of overly biased partitions. The cyclic MinimaxSplit algorithm further ensures that each dimension is used in a balanced manner throughout the tree construction process, avoiding dominance by a subset of features.
A key theoretical result of this study is that the cyclic MinimaxSplit algorithm achieves an exponential MSE decay rate of given sufficient data samples, without requiring additional variance decay assumptions (Theorem 9). Furthermore, we derived empirical risk bounds for this method, establishing its robustness in different settings (Theorem 10). Along the same lines of techniques, we prove novel results on the convergence rates of univariate partition-based martingale approximations, which are of their own interest.
In addition to single-tree regression analysis, we explored ensemble learning approaches, where we combined decision trees constructed using MinimaxSplit, cyclic MinimaxSplit, and VarianceSplit methods within a random forest framework. The results demonstrated that a hybrid approach leveraging different splitting techniques yields superior performance, particularly when dealing with non-uniform data distributions (Section 3.4.4). These findings suggest that the MinimaxSplit approach can provide more stable and adaptive decision trees compared to traditional methods, especially in high-dimensional settings.
Future research could explore the MinimaxSplit algorithm in classification tasks, particularly for imbalanced datasets. We expect that an adaptive splitting strategy could replace cyclic selection, dynamically adjusting to feature importance. These directions may refine the theoretical foundations of MinimaxSplit and broaden its practical applications in high-dimensional and complex learning environments.
Acknowledgment
The authors thank Ruodu Wang for his valuable input. We provide our reproducible code at github.com/hrluo. HL was supported by U.S. Department of Energy under Contract DE-AC02-05CH11231 and U.S. National Science Foundation NSF-DMS 2412403.
Appendix A Proofs and further results
A.1 Proof of Theorem 3
Variance martingale
@@ A.1 Proof of Theorem 3
The first observation is that the split location is precisely the middle point of the two means of the two descendants. This is a special case of Theorem 1 of Ishwaran (2015) applied with $f$ monotone and $X\sim U(0,1)$, but for completeness, we provide a full argument here with our notation. To see this, consider (without loss of generality by shifting) the law of $U$ with mean zero, and let $u=u_{*}$ be the optimal split location. Let $p_{1},p_{2}$ and $m_{1},m_{2}$ be the associated one-step probabilities and locations in the binary tree representation. Since the second moment $\mathbb{E}[M_{1}^{2}]$ is maximized, we must have
$0=\frac{\mathrm{d}}{\mathrm{d}u}\mathbb{E}[M_{1}^{2}]\mid_{u_{*}}=\frac{\mathrm{d}}{\mathrm{d}u}(p_{1}m_{1}^{2}+p_{2}m_{2}^{2})\mid_{u_{*}}.$ (32)
Observe that
$\frac{\mathrm{d}p_{1}}{\mathrm{d}u}=-\frac{\mathrm{d}p_{2}}{\mathrm{d}u}=\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}$
and
$\frac{\mathrm{d}m_{1}}{\mathrm{d}u}\mid_{u_{*}}=\frac{(u_{*}-m_{1})}{p_{1}}\,\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}\mid_{u_{*}};\leavevmode\nobreak\ \frac{\mathrm{d}m_{2}}{\mathrm{d}u}\mid_{u_{*}}=\frac{(m_{2}-u_{*})}{p_{2}}\,\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}\mid_{u_{*}}.$
Inserting into (32) leads to
$\frac{\mathrm{d}\mathbb{P}_{U}}{\mathrm{d}u}\mid_{u_{*}}\Big{(}m_{1}^{2}-m_{2}^{2}+2m_{1}(u_{*}-m_{1})-2m_{2}(u_{*}-m_{2})\Big{)}=0,$
which simplifies into $m_{1}+m_{2}=2u_{*}$ (in the case $\mathrm{d}\mathbb{P}_{U}/\mathrm{d}u(u_{*})=0$, we replace $u_{*}$ with the closest point of $u_{*}$ to $\operatorname{supp}U$ and this does not change the values of $p_{1},p_{2},m_{1},m_{2}$). This fact will be frequently used in the following.
At level $k$, $\mathbb{E}[(M_{k-1}-M_{k})^{2}]$ is a sum of $2^{k}$ terms in the binary tree representation (10), which we may rewrite as $\sum_{j}p_{j}d_{j}^{2}$ using a change of variable. This corresponds to the $2^{k}$ edges (indexed by $[2^{k}]$) in the binary tree between levels $k-1$ and $k$, and $p_{j}$ represents the probabilities of the edge $j$ and $d_{j}$ is the location difference between the end vertices of the edge, or the length of the edge. Our goal is to bound from above
@@
That is, the variance increases by splitting the $j$-th mass at level $k-1$ (which is the common ancestor of the $(2j-1)$-th and $(2j)$-th masses). By the martingale property,
@@
The general strategy to analyze the quantity (33) is to start from an arbitrary mass/node $\mathbf{M}$ at level $k-1$, trace back its ancestors, and bound from above the probability of the mass (which is $p_{2j-1}+p_{2j}$ in the above expression) and the upper bounds for the length of the edges (which is $d_{2j-1},\,d_{2j}$).
The path leading to a mass $\mathbf{M}$ at level $k-1$ in the binary tree representation (10) consists of $k$ nodes: $\emptyset=\mathbf{N}_{0},\mathbf{N}_{1},\dots,\mathbf{N}_{k-1}=\mathbf{M}$, where $\mathbf{N}_{j}\in V_{j}$. We denote the lengths of the edges connecting the node $\mathbf{N}_{j-1}$ to its two descendants on level $j$ by $a_{j},b_{j},\,j\in[k-1]$, where $b_{j}=|\ell_{\mathbf{N}_{j-1}}-\ell_{\mathbf{N}_{j}}|$. It follows by the martingale property that the mass $\mathbf{M}$ has probability
@@
Let $a_{0}=b_{0}=1\geqslant\sup U-\inf U$. We consider the following algorithm, which starts from the tuple $(a_{0},b_{0},a_{1},b_{1},q_{1})$ and updates it $k-1$ times for each $j\in[k-1]$ when we trace the path down to $\mathbf{M}$. At step $j$, the tuple represents the two pairs $(a_{k},b_{k}),\,(a_{j},b_{j})$ that govern the maximum length of the edges of the $j$-th node (that is, the distance from its location to the boundaries of the intervals this node represents) for a certain $j\in[k-1]$, as well as the probability weight $q_{j}$ multiplied at this step. In a generic setting, suppose that we have $(a_{k},b_{k},a_{j},b_{j},q_{j})$ updated at the node $\mathbf{N}_{j}\in V_{j}$. This means that the edges forming $a_{k+1},\dots,a_{j}$ are all on one side of the path since only the last one of these contributes to the bound. Now when updating the node at level $j+1$ there are two cases:
@@
Here, we have used (34) and the fact that $m_{1}+m_{2}=2x_{*}$ where $x_{*}$ is the split location.
Let $(a,b,a^{\prime},b^{\prime},q)$ be generic entries of the tuple and $(\hat{a},\hat{b},\hat{a}^{\prime},\hat{b}^{\prime},\hat{q})$ be the updated tuple, where we recall $\hat{q}=\hat{a}^{\prime}/(\hat{a}^{\prime}+\hat{b}^{\prime})$. We claim that
@@
This proves the claim.
Recall that our goal is to bound (33). The probability of the mass at each step is multiplied by $q$, and the edge lengths are bounded by $(a+b)/2$ and $(a^{\prime}+b^{\prime})/2$. By the above claim, their product multiplies by at most $1/2$ each time we update the tuple. We have thus shown that
@@
[…]
@@
Let $\phi(u)=\mathbb{P}(U<u)\mathrm{Var}(U\mid U<u)$ and $\psi(u)=\mathbb{P}(U\geqslant u)\mathrm{Var}(U\mid U\geqslant u)$. Note that $\phi$ is increasing and $\psi$ is decreasing in $u$, and both functions are continuous if we assume that $U$ is atomless. Recall by the minimax property that at each step we pick a $u\in\mathbb{R}$ such that $\max\{\phi(u),\psi(u)\}$ is minimized. This means that $\phi(u)=\psi(u)$. Since $\mathrm{Var}(U)\geqslant\phi(u)+\psi(u)$, we have both $\phi(u)\leqslant\mathrm{Var}(U)/2$ and $\psi(u)\leqslant\mathrm{Var}(U)/2$. Inductively, we have for any $k$, $\max(p_{j}d_{j}^{2})\leqslant 2^{-k}\mathrm{Var}(U)\leqslant 2^{-2-k}$. Combining the above, we arrive at

### v2

@@ A.4 Proof of Theorem 16
If the support of $U$ is a connected interval, Lemma A.1 is a special case of Theorem 1 of Ishwaran (2015) applied with $f$ monotone and $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$.
@@
Recall the binary subtree representation (17). Since $U$ is atomless, the binary subtree is in fact always a binary tree. At level $k$, $\mathbb{E}[(M_{k-1}-M_{k})^{2}]$ is a sum of $2^{k}$ terms, which we may rewrite as $\sum_{j}p_{j}d_{j}^{2}$ using a change of variable. This corresponds to the $2^{k}$ edges (indexed by $j\in[2^{k}]$) in the binary tree between levels k1k-1 and $k$, and $p_{j}$ represents the probabilities of the edge $j$ and $d_{j}$ is the location difference between the end vertices of the edge, or the length of the edge. Our goal is to bound from above
@@
That is, the variance increases by splitting the $j$-th mass at level k1k-1 (which is the common ancestor of the (2j1)(2j-1)-th and $(2j)$-th masses). By the martingale property,
@@
The general strategy to analyze the quantity (37) is to start from an arbitrary mass/node $\mathbf{M}$ at level k1k-1, trace back its ancestors, bound from above the probability of the mass (which is $p_{2j-1}+p_{2j}$ in the above expression), and establish upper bounds for the length of the edges (which is $d_{2j-1},\,d_{2j}$).
The path leading to a mass $\mathbf{M}$ at level k1k-1 in the binary tree representation (17) consists of $k$ nodes: $\emptyset=\mathbf{N}_{0},\mathbf{N}_{1},\dots,\mathbf{N}_{k-1}=\mathbf{M}$, where $\mathbf{N}_{j}\in V_{j}$. We denote the lengths of the edges connecting the node $\mathbf{N}_{j-1}$ to its two descendants on level $j$ by $a_{j},b_{j},\,j\in[k-1]$, where $b_{j}=|\ell_{\mathbf{N}_{j-1}}-\ell_{\mathbf{N}_{j}}|$. It follows by the martingale property that the mass $\mathbf{M}$ has probability
@@
Let $a_{0}=b_{0}=1\geqslant\sup U-\inf U$. We consider the following algorithm, which starts from the tuple $(a_{0},b_{0},a_{1},b_{1},q_{1})$ and updates it k1k-1 times for each $j\in[k-1]$ when we trace the path down to $\mathbf{M}$. At step $j$, the tuple represents the two pairs $(a_{k},b_{k}),\,(a_{j},b_{j})$ that govern the maximum length of the edges of the $j$-th node (that is, the distance from its location to the boundaries of the intervals this node represents) for a certain $j\in[k-1]$, as well as the probability weight $q_{j}$ multiplied at this step. In a generic setting, suppose that we have $(a_{k},b_{k},a_{j},b_{j},q_{j})$ updated at the node $\mathbf{N}_{j}\in V_{j}$. This means that the edges forming $a_{k+1},\dots,a_{j}$ are all on one side of the path since only the last one of these contributes to the bound. Now when updating the node at level $j+1$ there are two cases:
@@
Here, we have used (38) and Lemma A.1.
[…]
@@
This proves (39).
Recall that our goal is to bound (37). The probability of the mass at each step is multiplied by $q$, and the edge lengths are bounded by $(a+b)/2$ and $(a^{\prime}+b^{\prime})/2$. By (39), their product multiplies by at most $1/2$ each time we update the tuple. We have thus shown that
@@
It is noted in Zhang et al. (2024) that the Simons martingale satisfies the separated tree condition (i.e., the branches of the tree representation from all different levels, when projected onto the real line as intervals, are either disjoint or have a containment relationship), which is a consequence of the construction. We follow a similar argument to the variance martingale. Consider a node $\mathbf{M}$ at level k1k-1 and a path $(\emptyset=\mathbf{N}_{0},\dots,\mathbf{N}_{k-1}=\mathbf{M})$ leading to $\mathbf{M}$, where $\emptyset$ denotes the root of a binary tree. We provide an algorithm that recursively defines the tuples $(A_{j},B_{j},q_{j}),\leavevmode\nobreak\ 0\leqslant j\leqslant k-1$ for each node $\mathbf{N}_{j},\leavevmode\nobreak\ 0\leqslant j\leqslant k-1$ on that path, from the root to $\mathbf{M}$. The quantities $A_{j}$ and $B_{j}$ represent the maximum possible lengths of the two edges emanating from a node $\mathbf{N}_{j}$, and $q$ represents the one-step splitting probability leading to $\mathbf{N}_{j}$, based on the constraints from the separated tree property.
@@
Let $\phi(u)=\mathbb{P}(U<u)\mathrm{Var}(U\mid U<u)$ and $\psi(u)=\mathbb{P}(U\geqslant u)\mathrm{Var}(U\mid U\geqslant u)$. Note that $\phi$ is increasing and $\psi$ is decreasing in $u$, and both functions are continuous if we assume that $U$ is atomless. Recall by the minimax property that at each step we pick $u\in\mathbb{R}$ such that $\max\{\phi(u),\psi(u)\}$ is minimized. This means that $\phi(u)=\psi(u)$. Since $\mathrm{Var}(U)\geqslant\phi(u)+\psi(u)$, we have both $\phi(u)\leqslant\mathrm{Var}(U)/2$ and $\psi(u)\leqslant\mathrm{Var}(U)/2$. Inductively, we have for any $k$, $\max(p_{j}d_{j}^{2})\leqslant 2^{-k}\mathrm{Var}(U)\leqslant 2^{-2-k}$. Combining the above, we arrive at